A sailboat tacking against the wind moves as follows: 4.2 km at 33° east of north, 3.6 km at 26° west of north, and 3.8 km at 41° east of north. The entire trip takes 1 h and 25 min. What is the displacement of the sailboat at the end of the trip?


1
Expert's answer
2021-11-07T19:26:23-0500

"x=4.2\cdot\sin33°-3.6\cdot\sin26°+3.8\cdot\sin41°=3.2\ (km)"


"y=4.2\cdot\cos33°+3.6\cdot\cos26°+3.8\cdot\cos41°=9.6\ (km)"


"r=\sqrt{3.2^2+9.6^2}=10.1\ (km)" . Answer